Solve




x ^ 2 * (d ^ 2 * y)/(d * x ^ 2) - 2x * (dy)/(dx) - 4y = x ^ 2 + 2 * log x

1
Expert's answer
2021-11-30T13:59:24-0500
"x^2y''-2xy'-4y=x^2+2\ln x"

The corresponding homogeneous differental equation


"x^2y''-2xy'-4y=0"

Auxiliary equation


"x^2(x^r)''-2x(x^r)'-4x^r=0"

"x^r(r(r-1)-2r-4)=0"

"r^2-3r-4=0"

"(r+1)(r-4)=0"

"r_1=-1, r_2=4"

The general solution of the homogeneous differential equation is


"y_h=\dfrac{c_1}{x}+c_2x^4"

Find the particular solution of the non-homogeneous differential equation


"y_p=Ax^2+Bx+C+D\ln x"

"y_p'=2Ax+B+\dfrac{D}{x}"

"y_p''=2A-\dfrac{D}{x^2}"

Substitute


"2Ax^2-D-4Ax^2-2Bx-2D"

"-4Ax^2-4Bx-4C-4D\ln x=x^2+2\ln x"

"-6A=1"

"-6B=0"

"-3D-4C=0"

"D=-\dfrac{1}{2}"

"A=-\dfrac{1}{6}, B=0, C=\dfrac{3}{8}"

"y_p=-\dfrac{1}{6}x^2+\dfrac{3}{8}-\dfrac{1}{2}\ln x"

The general solution of the given non-homogeneous differential equation is


"y=\dfrac{c_1}{x}+c_2x^4-\dfrac{1}{6}x^2+\dfrac{3}{8}-\dfrac{1}{2}\ln x"